OpenStudy (anonymous):

If independent random variable X and Y have variances 36 and 16 resp. What should be the correlation coefficient between (X+Y) and (X-Y) ?

OpenStudy (anonymous):

@amistre64 any idea?

OpenStudy (amistre64):

well, the respect variances would just be the addition or subtraction of the parts right? we aint go not -1Y or 4X stuff to get in the way

OpenStudy (amistre64):

how does variance relate to a correlation coeff?

OpenStudy (anonymous):

as per my knowledge correlation coeff is given by \[r(x,y)=\frac{ Covariance of x and y }{ \sigma x*\sigma y }\]

OpenStudy (anonymous):

here we have to fine for x+y and x-y instead of x and y

OpenStudy (amistre64):

since X and Y are independant; VAR(X+Y) = VAR(X) + VAR(Y)

OpenStudy (amistre64):

but all the covariance stuff i see has to do with the expected value E(XY) = E((X-Xu)(Y-Yu))

OpenStudy (anonymous):

yeah that's right variance of x+y=52 and variance of x-y=20

OpenStudy (anonymous):

yeah that's where I am stuck how to find covariance if only variances are known

OpenStudy (amistre64):

i read that the covar of independant events is zero ....

OpenStudy (amistre64):

but this is the correlation between x-y and x+y .... not between x and y

OpenStudy (anonymous):

yep x and y are independent but x+y and x-y are dependent hence there should be covariance between them

OpenStudy (amistre64):

so at least the under part is known :) sqrt(52*20)

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

you dont have parts of this buried in another problem do you? sometimes they say, use this from that in here to do stuff with

OpenStudy (anonymous):

nope nothing like that. I have posted the question as appeared in the university question paper. :(

OpenStudy (amistre64):

i recall reading something about this a few months back ... but cant quite find the site i saw it on

OpenStudy (anonymous):

okay that's fine just let me know if you think of anything.

OpenStudy (amistre64):

one thing i have rattling around up in my head is: if we assume X to have a mean of 0; then mean X+Y = Yu mean X-Y = -Yu not too sure how to use that tho

OpenStudy (anonymous):

why would it have mean zero?

OpenStudy (amistre64):

dunno, its just something in my head ..... lets say we have a set of number a,b,c,....,k whose mean is m, and var is q an equivalent set of a-m,b-m,c-m,...,k-m will have a mean of 0, and retain the same var of q

OpenStudy (amistre64):

the relationship between X and Y is the same regardless if we move at all such that X has a mean of 0

OpenStudy (amistre64):

without knowing a mean for the Y, that simply leaves Y as some unknown

OpenStudy (anonymous):

okay lets assume but still how we can find E(XY) = E((X-Xu)(Y-Yu))

OpenStudy (amistre64):

i believe we want \[E(X-Y)(X+Y)\] \[E([X-Y]+Yu)([X+Y]-Yu)\]

OpenStudy (amistre64):

not sure how this would help tho

OpenStudy (anonymous):

yeah that's exactly what we want....

OpenStudy (amistre64):

can we construct that from using standard deviations as random X and Y values?

OpenStudy (anonymous):

ahh don't think so. can't remember any formula for such condition.

OpenStudy (amistre64):

say: X = {-18, -12, -6,0, 6, 12, 18} Y = {-12,-8,-4, 0,4,8,12}+Yu

OpenStudy (anonymous):

okay then?

OpenStudy (amistre64):

then use those to determine X+Y and X-Y and run those values thru Sxy Sxx and Syy

OpenStudy (amistre64):

in other words, start from scratch to see if we can get to a viable result Excel might help

OpenStudy (anonymous):

okay I'll do it later now I got to go..

OpenStudy (amistre64):

well, heres what i got wroted up, might want to chk it for mistakes tho :)

OpenStudy (amistre64):

after a got home last night, i realized why i was centering X at zero, and for that matter Y as well. Notice that the operation: X-Xu does this by default. take Xu - Xu, = 0, so Xu is shifted to 0, and all the X values about Xu are shifted center around 0 as well X-Xu is X "centered at 0" Y-Yu is Y "centered at 0"

OpenStudy (amistre64):

if we take Xu and Yu to be some fixed interval apart and move the whole shebang to X = 0, we retain that same relationship regardless of where they actually started at. :